OpenStudy (anonymous):

Suppose you are solving the system of equations below using the substitution method. You solve for $ y $ in the first equation and substitute it into the second equation. What equation must you solve then? 4x+y=5 x^2+9y=0

OpenStudy (anonymous):

i presume you are trying to solve 4x + y = 5 ----(eqn 1) x^2 + 9y = 0 ---(eqn 2) rearranging (1): y = 5 - 4x --- (eqn 3) sub (3) into (2): x^2 + 9 (5-4x) = o x^2 - 36x + 45 = 0 solve the quadratic you'll get your x and sub it back in to another eqn to find your y.

OpenStudy (anonymous):

y = -4x +5 sub in second x^2 + 9 ( -4x + 5 ) = 0 x^2 -36x +45 = 0 the answer is so ugly, copy it from here http://www.wolframalpha.com/input/?i=4x%2By%3D5+and+x^2%2B9y%3D0